# Energy-Efficient Designs for SIM-Based Broadcast MIMO Systems

### v1

@@ VIII Conclusion
In this paper, we studied the EE maximization in a SIM-aided broadcast system with DPC and linear precoding at the BS. For DPC, we exploited the well-known BC-MAC duality and optimize the users’ covariance matrices by employing a SCA-based technique, which establishes a tight lower bound of the achievable sum-rate, and applying Dinkelbach’s method. A similar approach was used to optimize the precoders in the case of linear precoding. The phase shifts of the SIM meta-elements for DPC and linear precoding were optimized using a conventional projected gradient-based method due to its simplicity. Also, we conducted a computation complexity analysis of the proposed optimization algorithms and proved their convergence. Numerical results showed that implementing these proposed optimization algorithms can significantly improve the EE for SIM-aided broadcast systems. Moreover, we demonstrated that the EE depends on the number of SIM meta-elements and their distribution across the SIM layers. We also found that in SIM-aided broadcast systems without precoding, optimal energy efficient transmission strategies typically involve a subset of active transmit antennas.
Appendix A Proof of Theorem 2
@@
$\mathbf{G}_{2}=$ $\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{2}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}$ (72)
@@
From this gradient expression for the achievable rate of user $k$, we can easily obtain the appropriate gradients for all other users. After summation of all these gradient expressions, we obtain (58). This completes the proof.

### v2

@@ VIII Conclusion, Insights, and Future Work
In this paper, we studied the EE maximization in a SIM-aided broadcast system, considering both DPC and LP at the BS. For DPC, we exploited the well-known BC-MAC duality to reformulate the optimization problem and employed a SCA-based technique to derive a tight lower bound on the achievable sum-rate, which was then optimized using Dinkelbach’s method. A similar approach was used to optimize precoders in the case of LP. The phase shifts of the SIM meta-elements were optimized using a conventional projected gradient-based method due to its simplicity. Also, we conducted a computation complexity analysis of the proposed optimization algorithms and proved their convergence.
We carried out extensive numerical experiments and the simulation results have lead to the following insights which are useful for designing energy-efficient SIM-based systems:
  • The proposed optimization algorithms significantly improve the EE for SIM-aided MIMO broadcast systems.
  • DPC outperforms LP in terms of EE due to its superior interference cancellation, but the EE improvement remains marginal, making LP a practical alternative with lower complexity.
  • SIM-based systems without digital precoding still suffer from higher multi-user interference, leading to lower EE.
  • Increasing the number of meta-elements improves EE, but with diminishing returns due to increased power consumption and the logarithmic nature of the achievable sum-rate.
  • Optimal meta-element distribution across layers is crucial; simply increasing the number of SIM layers can degrade performance due to inter-layer signal propagation loss.
  • Discrete phase shift quantization negatively impacts EE, especially for systems with more SIM layers.
  • In SIM-aided broadcast systems without digital precoding, optimal energy efficient transmission strategies typically involve a subset of active transmit antennas.
For future work, the above insights immediately motivate investigating the optimal distance between SIM layers to achieve the best EE performance. Another interesting topic is studying the performance of SIMs in a dynamic environment with user mobility, which would involve adapting the proposed optimization algorithms.
[Proof of Theorem 2]
@@
$\mathbf{G}_{2}=$ $\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{2}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}.$ (69)
@@
From this gradient expression for the achievable rate of user $k$, we can easily obtain the appropriate gradients for all other users. After summing all these gradient expressions, we obtain (55). This completes the proof.